# Stephanie: Step-by-Step Dialogues for Mimicking Human Interactions in Social Conversations

### v1

@@ 1 Introduction
In the field of natural language processing, the research and development of dialogue systems continue to advance. Some systems aim to mimic human communication in daily life [^1] https://nijigen.com.cn, [^2] https://chatgpt.com/?model=gpt-4o, providing a more natural user experience. However, these systems predominantly employ a Single-Step Dialogue Paradigm(Abbas et al., ; Touvron et al., 2023; Du et al., 2022; Abdin et al., 2024; Achiam et al., 2023), where the system provides a comprehensive, one-time response to each user input, quickly addressing user questions or needs. While this approach can deliver a wealth of information in a single response, it falls short in simulating the fluidity and emotional exchange of real human conversations. In reality, daily human conversations are ongoing, dynamically evolving processes involving multiple topics and emotional exchanges(Song et al., 2022; Butler, 2011; Nie et al., 2024; Poria et al., 2019), and the current single-step dialogue paradigm fails to adequately capture this complexity and richness of emotions.
@@
Figure 1: A single-step dialogue system and Stephanie. Stephanie constructs a dialogue composed of multiple dispersed yet coherent responses.
@@
To better emulate the style of human social conversations, this paper introduces an innovative dialogue paradigm named step-by-step dialogue (Stephanie), as shown in Figure 1. Unlike single-step dialogue, Stephanie mimics casual chats in instant messaging applications, creating a more natural and continuous dialogue flow with the power of in-context learning (Min et al., 2022; Chen et al., 2023a; Wang et al., 2024). Under this paradigm, the dialogue system does not just provide a one-time response to each input but constructs a conversation composed of multiple dispersed yet coherent responses. This design allows the system to gradually develop the conversation, with each response focusing on different aspects of the dialogue, making the conversation more detailed, rich, and emotionally nuanced. We found that it provides a better conversation experience, evoking greater user engagement. For example, in the step-by-step chat mode, the system can address various aspects of a user’s emotional expression step by step, first by supporting users through empathetic language and understanding, and then by asking questions or expanding the topic, gradually building deeper and more continuous emotional communication.
@@
If the one-time response of a single-step dialogue system is simply divided into multiple responses by punctuation, the overall logic and integrity of the one-time response itself will result in an unnatural and stiff step-by-step dialogue, which does not resemble a real social interaction with people. In order to fully consider the semantic similarities and differences between sentences, as well as naturalness and anthropomorphism when generating step-by-step dialogue and implementing a dialogue system with a step-by-step chat function, we introduced a comprehensive prompting framework that employs a dual learning strategy and a Further-Split post-editing method to generate and optimize step-by-step dialogue datasets. We then used this dataset with a specific fine-tuning strategy to be compatible with existing large models, thereby establishing a step-by-step dialogue system. The step-by-step dialogue paradigm demonstrates significant academic and practical value in enhancing the naturalness and emotional depth of chat systems. By simulating real social interactions, this research not only advances the technology of dialogue systems but also provides new insights and approaches for achieving more natural and human-like communication between machines and humans.
@@
  • We innovatively propose a step-by-step dialogue paradigm that utilizes a series of dispersed yet coherent responses to more closely mirror the style of real human communication interactions, thereby enhancing the emotional depth and human-likeness of the dialogue.
@@ 3.1 Dual Learning Strategy for Step-by-Step Dialogue Generation
  • Positive Learning Objectives: To help the model understand the step-by-step dialogue paradigm, we created five high-quality step-by-step dialogue examples as the negative objectives $P$. These examples simulate everyday social exchanges between two individuals and serve as a basis for few-shot learning, training the model to generate coherent and emotionally rich step-by-step dialogues in different background contexts.
  • Negative Learning Objectives: Simultaneously, we designed five high-quality single-step dialogue examples as the negative objectives $N$. Through contrastive learning, these examples enable the model to discern the differences between single-step and step-by-step dialogues. This negative learning approach helps the model better understand the step-by-step dialogue paradigm by pushing away dissimilar single-step dialogue examples.
@@ 3.3 Dataset Generation and Finetuning Strategy for Stephanie
During the finetuning process, we introduced delimiters to format the dataset, providing structured input and output for the model, where the content between each pair of delimiters represents a single exchange between the two dialogue participants. We then used this newly formatted step-by-step dialogue dataset to finetune the model. After finetuning, the model’s output also adopted the same delimiter-separated step-by-step dialogue format. Finally, we used a series of scripts and software tools to design a user-friendly UI interface that converts the input and output into a format similar to message bubbles in social software, allowing users to interact with the large language model using the step-by-step dialogue paradigm.

### v2

@@ 1 Introduction
In the field of natural language processing, the research and development of dialogue systems continue to advance. Some systems aim to mimic human communication in daily life [^1] https://nijigen.com.cn, [^2] https://chatgpt.com/?model=gpt-4o, providing a more natural user experience. However, these systems predominantly employ a Single-Step Dialogue Paradigm(Abbas et al., ; Touvron et al., 2023; Du et al., 2022; Abdin et al., 2024; Achiam et al., 2023), where the system provides a comprehensive, one-time response to each user input, quickly addressing user questions or needs. While this approach can deliver a wealth of information in a single response, it falls short in simulating the fluidity of real human conversations and attracting user engagement. In reality, daily human conversations are ongoing, dynamically evolving processes involving multiple topics (Song et al., 2022; Butler, 2011; Nie et al., 2024; Poria et al., 2019). However, the current Single-Step Dialogue Paradigm fails to fully capture this complexity and lacks in attracting users to engage in conversations.
@@
Figure 1: A single-step dialogue system and Stephanie. Stephanie constructs a dialogue composed of multiple dispersed yet coherent responses. ES stands for the engaging score given by humans, which is introduced in section 4.4.
@@
To better emulate the style of human social conversations, this paper introduces an innovative dialogue paradigm named step-by-step dialogue (Stephanie), where its human-rated engaging score surpasses that of single-step dialogue, as illustrated in Figure 1. Unlike single-step dialogue, Stephanie mimics casual chats in instant messaging applications, creating a more natural and continuous dialogue flow with the power of in-context learning (Min et al., 2022; Chen et al., 2023a; Wang et al., 2024). Under this paradigm, the dialogue system does not just provide a one-time response to each input but constructs a conversation composed of multiple dispersed yet coherent responses. This design allows the system to gradually develop the conversation, with each response focusing on different aspects of the dialogue, making the conversation more detailed, rich, and engaging. We found that it provides a better conversation experience, evoking greater user engagement. For example, in the step-by-step chat mode, the system can address various aspects of a user’s expression step by step, first by supporting users through empathetic language and understanding, and then by asking questions or expanding the topic, gradually building deeper and more continuous communication.
@@
If the one-time response of a single-step dialogue system is simply divided into multiple responses by punctuation, the overall logic and integrity of the one-time response itself will result in an unnatural and stiff step-by-step dialogue, which does not resemble a real social interaction with people. In order to fully consider the semantic similarities and differences between sentences, as well as naturalness and anthropomorphism when generating step-by-step dialogue and implementing a dialogue system with a step-by-step chat function, we introduced a comprehensive prompting framework that employs a dual learning strategy and a Further-Split post-editing method to generate and optimize step-by-step dialogue datasets. We then used this dataset with a specific fine-tuning strategy to be compatible with existing large models, thereby establishing a step-by-step dialogue system. The step-by-step dialogue paradigm demonstrates significant academic and practical value in enhancing the naturalness and engaging nature of chat systems. By simulating real social interactions, this research not only advances the technology of dialogue systems but also provides new insights and approaches for achieving more natural and human-like communication between machines and humans.
@@
  • We innovatively propose a step-by-step dialogue paradigm that utilizes a series of dispersed yet coherent responses to more closely mirror the style of real human communication interactions, thereby enhancing the engaging nature and human-likeness of the dialogue.
@@ 3.1 Dual Learning Strategy for Step-by-Step Dialogue Generation
  • Positive Learning Objectives: To help the model understand the step-by-step dialogue paradigm, we created five high-quality step-by-step dialogue examples as the positive objectives $P$. These examples simulate everyday social exchanges between two individuals and serve as a basis for few-shot learning, training the model to generate coherent and emotionally rich step-by-step dialogues in different background contexts.
  • Negative Learning Objectives: Simultaneously, we designed five high-quality single-step dialogue examples as the negative objectives $N$, which share the same theme as the five step-by-step dialogue examples for the positive objectives. Through contrastive learning, these examples enable the model to discern the differences between single-step and step-by-step dialogues. This negative learning approach helps the model better understand the step-by-step dialogue paradigm by pushing away dissimilar single-step dialogue examples.
@@ 3.3 Dataset Generation and Finetuning Strategy for Stephanie
During the finetuning process, we introduced delimiters to format the dataset, providing structured input and output for the model, where the content between each pair of delimiters represents a single exchange between the dialogue participants. We then used this newly formatted step-by-step dialogue dataset to finetune the model. After finetuning, the model’s output also adopted the same delimiter-separated step-by-step dialogue format. Then, we found Stephanie brings better user engagement, converting the delimiter format of the input and output into a format similar to message bubbles in social software, allowing users to interact with the large language model using the step-by-step dialogue paradigm.